rebecca has three coins in her pocket. they are all different from each other. tom has three coins and they are all the same as each other. tom has twice as much money as rebecca. what are the coins that each has?

Found 2 solutions by richwmiller, Edwin McCravy:
Answer by richwmiller(17219)

3t=2*r
r=a+b+c
2*(a+b+c)=3t
coins
1,5,10,25,50,100
if Tom has three alike he has
3,15,30,75,150 or 300
so Rebecca has a total of 1.5, 7.5, 15, 27.5, 75 or 150
she can't have 1.5, 7.5, 27.5
so she must have
15, 75 or 150
she needs three different coins to add up to 15, 75 or 150
5 and 10 add up to 15 but are only two coins
same with 75 cents 25 and 50 add up to 75
and also with 150
there are no current coins in circulation that meet the requirements



Answer by Edwin McCravy(20081)

This is not possible in the US because we do not have 20 cent coins here.

However it is possible in countries whose money unit is euros or pounds, 
EU or UK.

Rebecca has a 50 cents or pence coin, a 20 cents or pence coin, and a
5 cents or pence coin and has 75 cents or pence.

Tom has 3 50 cents or pence coins and has €1.50 or £1.50.

Edwin